OpenStudy (anonymous):

Office Expense. The monthly phone expense for an office is $1094 less than the janitorial expense. Three times the janitorial expense minus four times the phone expense is $248. What is the total of the two expenses?"

OpenStudy (anonymous):

Phone expense be x janitorial expense be y y - x = 1094 3y - 4x = 248

OpenStudy (anonymous):

y = 1094 + x 3 (1094 + x) - 4x = 248 3282 + 3x - 4x = 248 3282 -x = 248 x = 3034

OpenStudy (anonymous):

Thank you I am so lost I never should have went to college at 40 lol

OpenStudy (anonymous):

It's very brave to go to college at 40. Good job!

OpenStudy (anonymous):

Thank you sad think is my daughter is a wiz at math and is the 11th grade and says when she helps me I get frustrated when I do not understand and now she refuses to help uggg kids....

OpenStudy (anonymous):

I am Grade 10 and i am willing to help people who struggles in math.